OpenStudy (anonymous):

Please help? A segment 12" long is divided into two segments having lengths in the ratio of 2:3. Find the length of each segment. (Hint: Let 2x and 3x represent the length of the parts.) x = ? Segment 1 = (2x) = ? inches Segment 2 = (3x) = ? inches

OpenStudy (radar):

2x+3x=12 5x=12 x=12/5 2x=24/5" 3x=36/5" 60/5=12

OpenStudy (anonymous):

ohhhh okay it took me a bit to figure out how u got everything but i get it

OpenStudy (radar):

The two segments are 24/5 and 36/5 inches and they sum up to 12 inches.

OpenStudy (anonymous):

okay thanks a bunch cx

OpenStudy (radar):

you're welcome.